OpenStudy (anonymous):

Integral of arctan(8x) dx.

hartnn (hartnn):

put 8x=u then dx= (1/8)du ok?

OpenStudy (anonymous):

Alright.

hartnn (hartnn):

so your new integral is \[(1/8)\int\limits_{}^{}\tan^{-1}u \: du\] right ?

OpenStudy (anonymous):

Indeed it is.

hartnn (hartnn):

now u know the integral of tan^-1 x ??

OpenStudy (anonymous):

I actually don't have it memorized xD I just started integration by parts so we're doing hw problems to show the proofs for them.. so i'm assuming: dv = du v= u u= arctan(u) du= 1/1+u^2

hartnn (hartnn):

that is correct except, v=1

OpenStudy (anonymous):

i know it equals though: xarctan(x)-1/2ln(1+x^2)

OpenStudy (anonymous):

1, yea typo xD and i know what it is, but for some reason when i do the problem i don't get what it should.

hartnn (hartnn):

ok,what u got? how did u integrate u/(1+u^2) ?

hartnn (hartnn):

put another variable say y=1+u^2 dy = 2u du ok?

OpenStudy (anonymous):

yea, this is weird cuz usually i use u, but we kinda already used u-sub in the beginning xD but y works!

hartnn (hartnn):

ok,go ahead and ask if u still don't get it.

OpenStudy (anonymous):

alright i'll let ya know

OpenStudy (anonymous):

haha I got it xD i know what i was doing wrong too xD i kept writing ln(1+8x^2) instead of 64x^2 xD i'm a dunce.. thanks again hartnn :)

hartnn (hartnn):

welcome :)